OpenStudy (anonymous):

Solve the system of equations. 3x - 4y = -10 and 5x + 8y = -2 A. (1, 2) B. (2, 1) C. (1, -2) D. (-2, 1)

OpenStudy (anonymous):

Hi! What are your thoughts so far?

OpenStudy (anonymous):

(3x - 4y = -10)*2 5x + 8y = -2 6x-8y=-20 5x+8y=-2 ---------- 11x=-22 --->> x=-2 5(-2)+8y=-2 --->>> y=1 answer D